mommybear1821:

A copy machine makes 105 copies in 3 mins and 45 seconds. how many copies does it make per minute?

XLynx:

don't be Frash at math bruh 105/225 = x/60 x = 28 fahhhhhhhh

KyledaGreat:

To find how many copies the copy machine makes in one minute, we need to first convert the time of 3 minutes and 45 seconds to seconds. 3 minutes = 3 x 60 seconds = 180 seconds 45 seconds = 45 seconds Total time = 180 + 45 = 225 seconds Now we can find how many copies are made per second: 105 copies รท 225 seconds = 0.4667 copies per second To find how many copies per minute, we can multiply the result by 60 seconds: 0.4667 copies per second x 60 seconds = 28 copies per minute (rounded to the nearest whole number) Therefore, the copy machine makes approximately 28 copies per minute.